A test is given with 10 T-F questions. How many different ways can the answer sheet be filled out so that 8 or more questions are answered correctly? I am drawing a blank and do not know where to start?
Bob

Found 2 solutions by jorel1380, stanbon:
Answer by jorel1380(3719)

If you have 10 questions, of which 8 are solved correctly, then you have 10C8 different combinations of answers:
10C8=10x9x8!/2x1x8!=5x9=45 different combinations.

Answer by stanbon(75887)

A test is given with 10 T-F questions. How many different ways can the answer sheet be filled out so that 8 or more questions are answered correctly? I am drawing a blank and do not know where to start?
------
Binomial Problem with n = 8 and p = 1/2
---
P(8<= x <=10) = 1-P(0<= x <=7)
= 1 - binomcdf(8,0.5,7)
= 0.0039
============
Cheers,
Stan H.